[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri May 13, 2016

TSB Admin nhc.tsbadmin at noaa.gov
Fri May 13 13:20:23 UTC 2016


Here is today's summary of NHC computer operations:
--- Mark

----------

TAFB:
The Meteosat data was restored late yesterday. The problem was with a  
router at Wallops.

Jessica reported the following: The surface pressures in the Pacific for  
the surface progs on N-AWIPS came out as 5 digit numbers with tripled over  
isobars. Everything looked normal when we just viewed the pressures in  
N-AWIPS from GFE. Everything looked normal in GFE. However, there has been  
a change to combine periods, making the surface progs a grid valid 12 hours  
instead of 6 hours, since I worked last. Matt Sardi will investigate.
HSU:
Mike Brennan worked with TSB to update the support contact list. The  
updated versions will be posted in HSU and TAFB.
Other items:
The AWIPS load test yesterday was a success. There were no significant  
issues with doing the storm surge products and OPC back-up on the same day.  
Raytheon will be sending us a detailed analysis of the system usage for the  
test period (12-18 UTC) to see how close to capacity we were.
NHC RFCs:
None

-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160513/c7e4e079/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list